Let N be a random variable with PMF P(N Also XINnBinomial (n, 0) i) Pi, i-1, 2, il pi -1,0 Pi 1. I. Show that (X,N) is a minimal sufficient statistic for θ. 2. Show that is unbiased for θ.
Answer #1

1. A statistic is said to be minimal sufficient statistic for the parameter heta if it satisfies the following conditions:

a) if it is a sufficient for the parameter heta

b) if it is a function of every other sufficient statistic

The pmf of X|N=n is given by :

P[X|N=n] = (nCx )*hetax(1-heta)n-x , x=0(1)n

= 0 , otherwise

Since we have a single random variable as our sample we can consider it to be our statistic

Using Neyman Fisher Factorisation Theorem, we can say that since we can write:

f(x,n|heta) = (nCx )*hetax(1-heta)n-x

  = h(x).g(x,n ; heta)

we can say that (X,N) are our sufficient statistic for heta.

Then we know that any further statistic of heta will always be a function of X, we can say that X is the minimal sufficient statistic for heta.

2. Since we know that X|N=n ~ Bin(n,heta), we know that

0 h

に1 x-0

EGI-Σ Σ 즈-3c.PL1-. θ)-x.pi に1 x-0

small E[rac{X}{N}]=(sum_{i=1}^{infty }p_{i})(sum_{x=0}^{n}rac{x}{n}._{x}^{n} extrm{C}. heta ^{x}(1- heta )^{n-x})

1

small E[rac{X}{N}]= heta ,for all values of heta

Hence X/N is an Unbiased Estimator of heta.